### v2

# Near-Optimal Approximation of Matrix Functions by Lanczos-FAN. Amsel, T. Chen, A. Greenbaum, C. Musco, and C. Musco

Abstract
We study the widely used Lanczos method for approximating the action of a matrix function $f(\mathbf{A})$ on a vector $\mathbf{b}$ (Lanczos-FA). For the function $f(\mathbf{A})=\mathbf{A}^{-1}$, it is known that, when $\mathbf{A}$ is positive definite, the $\mathbf{A}$-norm error of Lanczos-FA after $k$ iterations matches the optimal approximation from the Krylov subspace of degree $k$ generated by $\mathbf{A}$ and $\mathbf{b}$. In this work, we ask whether Lanczos-FA also obtains similarly strong optimality guarantees for other functions $f$.
We partially answer this question by showing that, up to a multiplicative approximation factor, Lanczos-FA also matches the optimal approximation from the Krylov subspace for any rational function with real poles outside the interval containing $\mathbf{A}$’s eigenvalues. The approximation factor depends on the degree of $f$’s denominator and the condition number of $\mathbf{A}$, but not on the number of iterations $k$. This result provides theoretical justification for the excellent performance of Lanczos-FA on functions that are well approximated by rationals. Additionally, using different techniques, we prove that Lanczos-FA achieves a weaker notion of optimality for the functions $\mathbf{A}^{1/2}$ and $\mathbf{A}^{-1/2}$. Experiments confirm that our new bounds more accurately predict the convergence of Lanczos-FA than existing bounds, in particular those based on uniform polynomial approximation.
keywords:
matrix function approximation, Lanczos, Krylov subspace method, optimal approximation, rational functions
{MSCcodes}
65F60, 65F50, 68Q25
1 Introduction
Given a symmetric matrix $\mathbf{A}\in\mathbb{R}^{d\times d}$, the Lanczos method for matrix function approximation (Lanczos-FA) is perhaps the most widely used method for approximating the action of a matrix function $f(\mathbf{A})$ on a vector $\mathbf{b}\in\mathbb{R}^{d}$. Here, $f:\mathbb{R}\rightarrow\mathbb{R}$ is a scalar function and $f(\mathbf{A})$ is defined as $f(\mathbf{A}):=\sum_{i=1}^{d}f(\lambda_{i})\mathbf{u}_{i}\mathbf{u}_{i}^{\mathsf{T}}$, where $\mathbf{A}$ has eigendecomposition $\mathbf{A}=\sum_{i=1}^{d}\lambda_{i}\mathbf{u}_{i}\mathbf{u}_{i}^{\mathsf{T}}$. Like other Krylov subspace methods, Lanczos-FA iteratively constructs an approximation to $f(\mathbf{A})\mathbf{b}$ from the Krylov subspace
$\mathcal{K}_{k}(\mathbf{A},\mathbf{b}):=\operatorname{span}\{\mathbf{b},\mathbf{A}\mathbf{b},\ldots,\mathbf{A}^{k-1}\mathbf{b}\}.$ (1)
In particular, the $k$-th Lanczos-FA iterate is defined as
$\mathsf{lan}_{k}(f;\mathbf{A},\mathbf{b}):=\mathbf{Q}f(\mathbf{T})\mathbf{Q}^{\mathsf{T}}\mathbf{b},$ (2)
where $\mathbf{Q}=[\mathbf{q}_{1},\ldots,\mathbf{q}_{k}]$ is an orthonormal basis for the Krylov subspace $\mathcal{K}_{k}(\mathbf{A},\mathbf{b})$ such that $\operatorname{span}\{\mathbf{q}_{1},\ldots,\mathbf{q}_{j}\}=\mathcal{K}_{j}(\mathbf{A},\mathbf{b})$ for all $j\leq k$, and $\mathbf{T}:=\mathbf{Q}^{\mathsf{T}}\mathbf{A}\mathbf{Q}$. This algorithm can be implemented efficiently in $O(dk+k^{2}\log k)$ time, plus the time to compute $k$ matrix-vector products with $\mathbf{A}$; see for instance [MMS18, Appendix A].
Besides Lanczos-FA, a number of algorithms which construct approximations to $f(\mathbf{A})\mathbf{b}$ from the Krylov subspace $\mathcal{K}_{k}(\mathbf{A},\mathbf{b})$ have been developed [${}^{+}$02, ${}^{+}$20, CGMM23]. While many of these algorithms come with stronger theoretical guarantees than currently known for Lanczos-FA, their practical performance often seems to be worse (and rarely substantially better) than Lanczos-FA. Thus, despite being one of the oldest and simplest iterative methods for computing approximations to $f(\mathbf{A})\mathbf{b}$, Lanczos-FA remains popular in practice. The main goal of this paper is to improve our theoretical understanding of why the method performs so well.
1.1 Optimality guarantees for Krylov subspace methods
We are motivated by the remarkable observation that, for a wide variety of problem instances, Lanczos-FA converges almost as quickly as the best approximation to $f(\mathbf{A})\mathbf{b}$ that could be returned by any Krylov subspace method run for the same number of iterations. In particular, all Krylov subspace methods output approximations that lie in the span of $\mathcal{K}_{k}(\mathbf{A},\mathbf{b})$, that is, approximations of the form $p(\mathbf{A})\mathbf{b}$ for a polynomial $p$ of degree less than $k$. Given a problem instance $(f,\mathbf{A},\mathbf{b},k)$, the best possible approximation returned by a Krylov method is [^1] We choose to measure error in the Euclidean norm, and thus define optimality with respect to this norm. While other norms have been considered in prior work on matrix-function approximation [CGMM23], the Euclidean norm is the simplest and most natural starting point.
$\mathsf{opt}_{k}(f;\mathbf{A},\mathbf{b}):=\operatornamewithlimits{argmin}_{\mathbf{x}\in\mathcal{K}_{k}(\mathbf{A},\mathbf{b})}\|f(\mathbf{A})\mathbf{b}-\mathbf{x}\|_{2}.$
By definition, the error of this Krylov optimal iterate can be characterized as follows:
$\|f(\mathbf{A})\mathbf{b}-\mathsf{opt}_{k}(f;\mathbf{A},\mathbf{b})\|_{2}=\min_{\deg(p)<k}\|f(\mathbf{A})\mathbf{b}-p(\mathbf{A})\mathbf{b}\|_{2}.$ (3)
While, in general, no efficient algorithm for computing $\mathsf{opt}_{k}(f;\mathbf{A},\mathbf{b})$ is known, as shown in Fig. 1, the solution returned by Lanczos-FA often nearly matches the error of this best approximation. We are motivated to ask whether, at least for some class of problems, Lanczos-FA satisfies the following strong notion of approximate optimality:
Definition 1.1 (Near Instance Optimality).
For an input instance $(f,\mathbf{A},\mathbf{b},k)$, we say that a Krylov method is near instance optimal with parameters $C,c$ if
$\|f(\mathbf{A})\mathbf{b}-\textsf{{alg}}_{k}(f;\mathbf{A},\mathbf{b})\|_{2}\leq C\min_{\deg(p)<ck}\|f(\mathbf{A})\mathbf{b}-p(\mathbf{A})\mathbf{b}\|_{2}.$
Above, $\textsf{{alg}}_{k}(f;\mathbf{A},\mathbf{b})$ denotes the output of the method obtained from the degree $k$ Krylov subspace $\mathcal{K}_{k}(\mathbf{A},\mathbf{b})$, i.e., the output after $k$ iterations.
Definition 1.1 is a near-optimality condition parameterized by two values, $C$ and $c$. Typically we think of $C$ as $\geq 1$ (which allows for the algorithm to return error slightly larger than the optimal approximation), and of $c$ as being $\leq 1$ (which allows us to compare against the best solution over a slightly smaller Krylov subspace). A near instance optimal guarantee is stronger when $C$ and $c$ are closer to $1$.
Notably, the right hand side in Definition 1.1 depends on the entire problem instance, $(f;\mathbf{A},\mathbf{b},k)$, which is why we call the guarantee “instance optimal”. The error of the Krylov optimal iterate for a given $f$ can also be upper bounded by a less precise quantity that is independent of $\mathbf{b}$ and only weakly dependent on $\mathbf{A}$. Specifically, if $\Lambda$ is the spectrum of $\mathbf{A}$ and $[\lambda_{\textup{min}},\lambda_{\textup{max}}]=\{\mathbf{x}^{\mathsf{T}}\mathbf{A}\mathbf{x}:\|\mathbf{x}\|_{2}=1\}$ is the field of values (FOV) of $\mathbf{A}$, we have:
$\|f(\mathbf{A})\mathbf{b}-\mathsf{opt}_{k}(f;\mathbf{A},\mathbf{b})\|_{2}$ $\leq\|\mathbf{b}\|_{2}\min_{\deg(p)<k}\left(\max_{x\in\Lambda}|f(x)-p(x)|\right)$ (4)
$\leq\|\mathbf{b}\|_{2}\min_{\deg(p)<k}\left(\max_{x\in[\lambda_{\textup{min}},\lambda_{\textup{max}}]}|f(x)-p(x)|\right).$ (5)
Figure 1: Lanczos-FA error $\|f(\mathbf{A})-\mathsf{lan}_{k}(f;\mathbf{A},\mathbf{b})\|_{2}$ at each iteration for several functions/spectra. “FOV Optimal” is the right hand side of Definition 1.3 and “Spectrum Optimal” is the right hand side of Definition 1.2, both with $C=c=1$. “Instance Optimal” is the right hand side of Definition 1.1 with $C=c=1$, which is a lower bound for all Krylov subspace methods, including Lanczos-FA. Lanczos-FA performs nearly instance optimally on a wide range of problems, far better than the bound of Fact 1 predicts. This is easily seen in the bottom plots, which show the ratio of the error of the Lanczos-FA iterate and the Krylov optimal iterate, $\mathsf{opt}_{k}(f;\mathbf{A},\mathbf{b})$.
These upper bounds on the error of $\mathsf{opt}_{k}(f;\mathbf{A},\mathbf{b})$ inspire natural alternative definitions of optimality that relax Definition 1.1.
Definition 1.2 (Near Spectrum Optimality).
For an input instance $(f,\mathbf{A},\mathbf{b},k)$, we say that a Krylov method is near spectrum optimal with parameters $C,c$ if
$\|f(\mathbf{A})\mathbf{b}-\textsf{{alg}}_{k}(f;\mathbf{A},\mathbf{b})\|_{2}\leq C\|\mathbf{b}\|_{2}\min_{\deg(p)<ck}\left(\max_{x\in\Lambda}|f(x)-p(x)|\right).$
Definition 1.3 (Near FOV Optimality).
For an input instance $(f,\mathbf{A},\mathbf{b},k)$, we say that a Krylov method is near FOV optimal with parameters $C,c$ if
$\|f(\mathbf{A})\mathbf{b}-\textsf{{alg}}_{k}(f;\mathbf{A},\mathbf{b})\|_{2}\leq C\|\mathbf{b}\|_{2}\min_{\deg(p)<ck}\left(\max_{x\in[\lambda_{\textup{min}},\lambda_{\textup{max}}]}|f(x)-p(x)|\right).$
Each type of optimality tightly relaxes the previous one. Definition 1.1 implies Definition 1.2 and, as we show in Subsection A.1, for any $f,k$ and $\mathbf{A}$, there exists a “worst-case” choice of $\mathbf{b}$ for which their bounds coincide; likewise, Definition 1.2 implies Definition 1.3, and for any $f$ and $k$ there exists a “worst-case” choice of $\mathbf{A}$ and $\mathbf{b}$ for which their bounds coincide. We also note that, with some additional assumptions on $\mathbf{b}$, spectrum optimality implies instance optimality.
Lemma 1.4.
Let $\{\mathbf{u}_{1},\ldots,\mathbf{u}_{d}\}$ be the eigenvectors of $\mathbf{A}$. Given a problem instance $(f,\mathbf{A},\mathbf{b},k)$, if an algorithm is nearly spectrum optimal with parameters $C$ and $c$, then it is nearly instance optimal with parameters $C\cdot\frac{\|\mathbf{b}\|_{2}}{\min_{j}|\mathbf{u}_{j}^{\mathsf{T}}\mathbf{b}|}$ and $c$.
A proof of Lemma 1.4 appears in Subsection A.2. The bound is useful e.g., when $\mathbf{b}$ is a random vector. For example, when $\mathbf{b}$ has independent and identically distributed Gaussian entries, $\frac{\|\mathbf{b}\|_{2}}{\min_{j}|\mathbf{u}_{j}^{\mathsf{T}}\mathbf{b}|}=O(d^{3/2})$ with high probability.
The three notions of near optimality given in Definition 1.1, Definition 1.2, and Definition 1.3, provide a clear and intuitive framework for understanding the convergence behavior of Krylov subspace algorithms like Lanczos-FA by comparing them to this optimal iterate. For example, one standard bound on the convergence of Lanczos-FA is the following well-known FOV optimality guarantee:
Fact 1.
For all problem instances $(f,\mathbf{A},\mathbf{b},k)$ where $\mathbf{A}$ is symmetric [^2] While we only study real symmetric matrices, the Crouzeix-Palencia Theorem can be used to generalize Fact 1 to arbitrary square matrices [CP17]. The domain of the maximization in Definition 1.3 should be changed from $[\lambda_{\textup{min}},\lambda_{\textup{max}}]$ to the field of values of $\mathbf{A}$, and the parameter $C$ should be changed from $2$ to $2(1+\sqrt{2})$., Lanczos-FA is nearly FOV optimal with parameters $C=2,c=1$.
Proof 1.5.
See [Saa92], [OSV12], or [MMS18, Theorem 4.1].
While powerful in its generality, as shown in Fig. 1, this bound is loose in many cases because it does not take into account properties of $\mathbf{A}$ such as isolated or clustered eigenvalues. Our goal in this paper is to improve this bound by proving spectrum and instance near optimality guarantees for various classes of functions $f$. We believe that studying Lanczos-FA through this lens is a promising way to further our understanding of its remarkable success in practice.
1.2 Existing near optimality guarantees for Lanczos-FA
With this goal in mind, we note that only a few instance optimality or spectrum optimality results are known for Lanczos-FA. Most directly, instance optimality trivially holds when $f$ is a polynomial of degree $<k$. In this case, it is well known that $\mathsf{lan}_{k}(f;\mathbf{A},\mathbf{b})$ exactly applies $f(\mathbf{A})\mathbf{b}$; i.e., $\|f(\mathbf{A})\mathbf{b}-\mathsf{lan}_{k}(f;\mathbf{A},\mathbf{b})\|_{2}=0$ [Saa92]. A more interesting special case is when $f(x)=1/x$ and $\mathbf{A}$ is positive definite. Then, the Lanczos-FA algorithm is mathematically equivalent to the well-known conjugate gradient algorithm for solving a system $\mathbf{A}\mathbf{x}=\mathbf{b}$. This implies that Lanczos-FA is the optimal approximation in the Krylov subspace with respect to the $\mathbf{A}$-norm; that is,
$\mathsf{lan}_{k}(1/x;\mathbf{A},\mathbf{b})=\tilde{p}(\mathbf{A})\mathbf{b},\qquad\tilde{p}:=\operatornamewithlimits{argmin}_{\deg(p)<k}\|\mathbf{A}^{-1}\mathbf{b}-p(\mathbf{A})\mathbf{b}\|_{\mathbf{A}}.$
This immediately yields near instance optimality in the Euclidean norm:
$\|\mathbf{A}^{-1}\mathbf{b}-\mathsf{lan}_{k}(1/x;\mathbf{A},\mathbf{b})\|_{2}\leq\sqrt{\kappa(\mathbf{A})}\min_{\deg(p)<k}\|\mathbf{A}^{-1}\mathbf{b}-p(\mathbf{A})\mathbf{b}\|_{2},$ (6)
where $\kappa(\mathbf{A})$ is the condition number $\lambda_{\textup{max}}/\lambda_{\textup{min}}$ of $\mathbf{A}$. That is, Definition 1.1 is satisfied with $C=\sqrt{\kappa(\mathbf{A})}$ and $c=1$.
Note that the best polynomial approximation to $1/x$ on $\mathcal{I}=[\lambda_{\textup{min}},\lambda_{\textup{max}}]$ already converges at a geometric rate [JM05]:
$\min_{\deg(p)<k}\max_{x\in\mathcal{I}}|1/x-p(x)|=\frac{8t^{k+1}}{(t^{2}-1)^{2}(\lambda_{\textup{max}}-\lambda_{\textup{min}})},\qquad t:=1-\frac{2}{1+\sqrt{\kappa(\mathbf{A})}}.$
That is, FOV optimality (Fact 1) suffices to prove the exponential convergence of the conjugate gradient method. However, it is well understood that conjugate gradient (and equivalently Lanczos-FA) often converge even faster, which is explained theoretically by the stronger instance and spectrum optimality guarantees that the methods satisfy[BK01, LS13, CLS22]. This illustrates the value of such guarantees over weaker bounds like Fact 1 for understanding the behavior of Lanczos-FA in general.
If $\mathbf{A}$ is not positive definite, $\mathbf{T}$ may have an eigenvalue at or near to zero and the error of the Lanczos-FA approximation to $\mathbf{A}^{-1}\mathbf{b}$ can be arbitrarily large. In fact, the same is true for any function $f$ which is much larger on $\mathcal{I}=[\lambda_{\textup{min}},\lambda_{\textup{max}}]$ than on $\Lambda$, the set of eigenvalues of $\mathbf{A}$. However, while the Lanczos-FA iterates may be bad at some iterations, the overall convergence of Lanczos-FA with $f(x)=1/x$ is actually good. In particular, the convergence of Lanczos-FA can be related to the MINRES algorithm [CG96], which produces an optimal approximation to $\mathbf{A}^{-1}\mathbf{b}$ with respect to the $\mathbf{A}^{2}$-norm. This allows certain optimality guarantees for MINRES to be transferred to Lanczos-FA, even on indefinite systems; see Subsection A.3 for details. In particular, for every $k$, there exists $k^{*}\leq k$ such that [^3] For indefinite matrices, $\kappa(\mathbf{A})$ denotes the ratio of the largest eigenvalue magnitude to the smallest eigenvalue magnitude.
$\|\mathbf{A}^{-1}\mathbf{b}-\mathsf{lan}_{k^{*}}(1/x;\mathbf{A},\mathbf{b})\|_{2}\leq\kappa(\mathbf{A})\left(\mathrm{e}\sqrt{k}+\frac{1}{\sqrt{k}}\right)\min_{\deg(p)<k}\|\mathbf{A}^{-1}\mathbf{b}-p(\mathbf{A})\mathbf{b}\|_{2}.$ (7)
While (7) does not quite fit the form of Definition 1.1 because of the $k^{*}$ on the left side, it is similar in spirit. Also note that the dependence on $k$ in the prefactor $\mathrm{e}\sqrt{k}+{1}/{\sqrt{k}}$ is not of great significance. Indeed, the convergence of the optimal polynomial approximation to $1/x$ on two intervals bounded away from zero is geometric.
Beyond the situations described above, the only other near optimality guarantee for Lanczos-FA that we are aware of is for the matrix exponential. In particular, [DGK98, (45)] shows that the Lanczos-FA iterate satisfies the guarantee
$\|\exp(-t\mathbf{A})\mathbf{b}-\mathsf{lan}_{k}(\exp(-tx);\mathbf{A},\mathbf{b})\|_{2}$
$\leq 3\|\mathbf{A}\|_{2}^{2}\>t^{2}\max_{0\leq s\leq t}\left(\min_{\deg(p)<k-2}\|\exp(-s\mathbf{A})\mathbf{b}-p(\mathbf{A})\mathbf{b}\|_{2}\right).$
Again, this bound does not quite fit Definition 1.1 due to the maximization over $s$. The authors of [DGK98] state, “It is not known whether the maximum over $s\in[0,t]$ in (45) can be omitted and $s$ set equal to $t$ in the right-hand side.”
1.3 Our contributions
In Section 2, we prove near instance optimality for a broad class of rational functions (Theorem 2.1). To the best of our knowledge, these are the first true instance optimality guarantees to be proven for any function besides $f(x)=1/x$. In Subsection 2.7, we discuss how results of this kind can be used to derive approximation guarantees for functions that are uniformly well approximated by rational functions. In Section 3, we prove spectrum optimality results for two especially important matrix functions: square root (Theorem 3.2) and inverse square root (Theorem 3.1). In Section 4, we present experimental evidence showing that, for many natural problem instances, our bounds are significantly sharper than the standard bound of Fact 1. We also comment on the tightness of our results with respect to their dependence on $f$ and the condition number $\kappa(\mathbf{A})$. We conclude with a discussion of next steps and open problems in Section 5.
2 Near-optimality for rational functions
In this section, we study the Lanczos-FA approximation to $r(\mathbf{A})\mathbf{b}$, where $r(x)$ is a rational function with real-valued poles that lie in $\mathbb{R}\setminus\mathcal{I}$, where $\mathcal{I}:=[\lambda_{\textup{min}},\lambda_{\textup{max}}]$. Specifically,
$r(x):=\frac{n(x)}{m(x)},$ (8)
where $n(x)$ is any degree $p$ polynomial and
$m(x)=(x-z_{1})(x-z_{2})\cdots(x-z_{q}),\qquad z_{i}\in\mathbb{R}\setminus\mathcal{I}.$
Since $z_{j}\not\in\mathcal{I}$, one of ±$\pm(\mathbf{A}-z_{j}\mathbf{I})$ is positive definite. For convenience, we define
$\mathbf{A}_{j}:=\begin{cases}+(\mathbf{A}-z_{j}\mathbf{I})&z_{j}<\lambda_{\textup{min}}\\ -(\mathbf{A}-z_{j}\mathbf{I})&z_{j}>\lambda_{\textup{max}}\\ \end{cases}.$ (9)
Our main result on rational functions is the following near instance optimality bound, which holds under a mild assumption on the number of iterations $k$:
Theorem 2.1.
Let $r(x)={n(x)}/{m(x)}$ be a degree $(p,q)$-rational function as in Eq. 8 and define $\mathbf{A}_{j}$ as in Eq. 9. Then, if $k>\max\{p,q-1\}$, the Lanczos-FA iterate satisfies the bound
$\|r(\mathbf{A})\mathbf{b}-\mathsf{lan}_{k}(r;\mathbf{A},\mathbf{b})\|_{2}\leq q\cdot\kappa(\mathbf{A}_{1})\cdot\kappa(\mathbf{A}_{2})\cdots\kappa(\mathbf{A}_{q})\min_{\mathrm{deg}(p)<k-q+1}\|r(\mathbf{A})\mathbf{b}-p(\mathbf{A})\mathbf{b}\|_{2}.$
Above, $\kappa(\mathbf{A}_{i})$ is the condition number of $\mathbf{A}_{i}$, the ratio of the largest to smallest magnitude eigenvalues. Theorem 2.1 shows that Lanczos-FA used to approximate $r(\mathbf{A})\mathbf{b}$ satisfies Definition 1.1 with
$C=q\cdot\kappa(\mathbf{A}_{1})\cdot\kappa(\mathbf{A}_{2})\cdots\kappa(\mathbf{A}_{q}),\qquad c=1-\frac{q-1}{k}.$
To the best of our knowledge, this is the first near-instance optimality guarantee for Lanczos-FA proven for a class of functions with multiple parameters. [^4] Disregarding the trivial case when $f$ is a polynomial of degree $<k$, and thus Lanczos-FA is exact (Lemma 2.2).
Importantly, while the prefactor $q\cdot\kappa(\mathbf{A}_{1})\cdot\kappa(\mathbf{A}_{2})\cdots\kappa(\mathbf{A}_{q})$ may be large, it does not depend on the number of iterations $k$. We also note that
$\kappa(\mathbf{A}_{1})\cdot\kappa(\mathbf{A}_{2})\cdots\kappa(\mathbf{A}_{q})\leq\gamma^{q},$
where
$\gamma:=\left(1+\frac{\lambda_{\textup{max}}-\lambda_{\textup{min}}}{\eta}\right),\qquad\eta:=\min_{j=1,\ldots,q}\min\{|z_{j}-\lambda_{\textup{min}}|,|z_{j}-\lambda_{\textup{max}}|\};$
i.e. $\eta$ is the minimum distance between the spectrum of $\mathbf{A}$ and the closest pole of $r(x)$. When $\mathbf{A}$ is positive-definite and the poles of $r(x)$ are non-positive, then $\gamma\leq\kappa(\mathbf{A})$.
As discussed in Section 5, we expect that Theorem 2.1 can be tightened by significantly reducing the prefactor. Nevertheless, as illustrated in Fig. 2, even in its current form, the bound improves on the standard bound of Fact 1 in many natural cases; i.e., it more tightly characterizes the observed convergence of Lanczos-FA. Finally, as discussed further in Subsection 2.7, we note that, beyond rational functions being an interesting function class in their own right, Theorem 2.1 has implications for understanding the convergence of Lanczos-FA for functions like $x^{-\alpha}$ that have much better uniform rational approximations than uniform polynomial approximations.
Figure 2: Despite its large prefactor, the bound of Theorem 2.1 qualitatively captures the convergence behavior of Lanczos-FA for rational functions. It can be tighter than the standard FOV optimality bound of Fact 1, even for a moderate number of iterations $k$. We use rational approximations to $\exp(-x/10)$ and $\log(x)$ for comparison with Fig. 1; see Section 4 for more details.
2.1 Proof Sketch
Our proof leverages the near-optimality of Lanczos-FA in computing $\mathbf{A}^{-1}\mathbf{b}$ to obtain a bound for general rational functions. For illustration, consider the simplest possible rational function for which we are unaware of any previous near-optimality bounds: $\mathbf{A}^{-2}\mathbf{b}$. The Lanczos-FA approximation for this function is $\mathbf{Q}\mathbf{T}^{-2}\mathbf{Q}^{\mathsf{T}}\mathbf{b}=\mathbf{Q}\mathbf{T}^{-1}\mathbf{Q}^{\mathsf{T}}\mathbf{Q}\mathbf{T}^{-1}\mathbf{Q}^{\mathsf{T}}\mathbf{b}$. Using triangle inequality and submultiplicativity, we can bound the error of the approximation as
$\|\mathbf{A}^{-2}\mathbf{b}-\mathsf{lan}_{k}(x^{-2};\mathbf{A},\mathbf{b})\|_{2}$
$\leq\|\mathbf{A}^{-2}\mathbf{b}-\mathbf{Q}\mathbf{T}^{-1}\mathbf{Q}^{\mathsf{T}}\mathbf{A}^{-1}\mathbf{b}\|_{2}+\|\mathbf{Q}\mathbf{T}^{-1}\mathbf{Q}^{\mathsf{T}}\mathbf{A}^{-1}\mathbf{b}-\mathbf{Q}\mathbf{T}^{-2}\mathbf{Q}^{\mathsf{T}}\mathbf{b}\|_{2}$
$\leq\|\mathbf{A}^{-2}\mathbf{b}-\mathbf{Q}\mathbf{T}^{-1}\mathbf{Q}^{\mathsf{T}}\mathbf{A}^{-1}\mathbf{b}\|_{2}+\|\mathbf{Q}\mathbf{T}^{-1}\mathbf{Q}^{\mathsf{T}}\|_{2}\cdot\|\mathbf{A}^{-1}\mathbf{b}-\mathbf{Q}\mathbf{T}^{-1}\mathbf{Q}^{\mathsf{T}}\mathbf{b}\|_{2}.$
The first term on the right hand side can be bounded because it is the error of the $\mathbf{A}$-norm optimal approximation to $\mathbf{A}^{-2}\mathbf{b}$ (indeed, $\mathbf{Q}\mathbf{T}^{-1}\mathbf{Q}^{\mathsf{T}}\mathbf{A}$ is the $\mathbf{A}$-norm projector onto Krylov subspace). The second term, on the other hand, is proportional to the error of the $\mathbf{A}$-norm optimal approximation to $\mathbf{A}^{-1}\mathbf{b}$; i.e. of the CG iterate. We can upper bound this optimal error for approximating $\mathbf{A}^{-1}\mathbf{b}$ in terms of the optimal error for approximating $\mathbf{A}^{-2}\mathbf{b}$ by using the fact that, if $p(x)$ is a good degree $k$ polynomial approximation to $x^{-2}$, then $xp(x)$ is a good approximation to $x^{-1}$.
Our proof of Theorem 2.1 generalizes the above approach. We write the Lanczos-FA error for approximating $r(\mathbf{A})\mathbf{b}$ in terms of the error of the optimal approximations to a set of simpler rational functions, and then reduce polynomial approximation of $r$ to polynomial approximation of each of these particular functions.
2.2 Notation
We first introduce notation used throughout. Given a rational function $r(x)=n(x)/m(x)$ with numerator degree $p$ denominator degree $q$, for $1\leq i\leq j\leq q$, we define:
$m_{i,j}(x):=\prod_{k=i}^{j}(x-z_{k}).$
We adopt the convention that $m_{j+1,j}(x):=1$. Note that $m_{i,j}$ is a $(j-i+1)$-degree polynomial. Define also
$r_{j}(x):=n(x)/m_{1,j}(x).$
Note that $m_{1,q}=m$, so $r_{q}=r(x)$. Recall that for any function $f$, the Lanczos-FA approximation to $f(\mathbf{A})\mathbf{b}$ is $\mathsf{lan}_{k}(f;\mathbf{A},\mathbf{b})=\mathbf{Q}f(\mathbf{T})\mathbf{Q}^{\mathsf{T}}\mathbf{b}$. Define
$\mathsf{err}_{k}(f):=f(\mathbf{A})\mathbf{b}-\mathsf{lan}_{k}(f;\mathbf{A},\mathbf{b}).$
Proving Theorem 2.1 amounts to proving an upper bound on $\|\mathsf{err}_{k}(r)\|_{2}$. Finally, for any symmetric positive definite matrix $\mathbf{M}$, define
$\mathsf{opt}_{k}(f,\mathbf{M};\mathbf{A},\mathbf{b}):=\operatornamewithlimits{argmin}_{\mathbf{x}\in\mathcal{K}_{k}(\mathbf{A},\mathbf{b})}\|f(\mathbf{A})\mathbf{b}-\mathbf{x}\|_{\mathbf{M}}.$
For brevity, in the analysis below we will write $\mathsf{lan}_{k}(f)$ and $\mathsf{opt}_{k}(f,\mathbf{M})$ in place of $\mathsf{lan}_{k}(f;\mathbf{A},\mathbf{b})$ and $\mathsf{opt}_{k}(f,\mathbf{M};\mathbf{A},\mathbf{b})$, since $\mathbf{A}$ and $\mathbf{b}$ are fixed throughout the analysis.
2.3 Simplifying $\mathsf{lan}_{k}(r_{j})$ and $\mathsf{opt}_{k}(r_{j},\mathbf{A}_{j})$
We begin with a few standard results that will be useful for the proof of Theorem 2.1.
Lemma 2.2 ([DK89, Saa92]).
For a polynomial $n(x)$, $\mathsf{lan}_{k}(n)=n(\mathbf{A})\mathbf{b}$ if $k>\deg(n)$.
Proof 2.3.
By definition of the Krylov subspace, $\mathbf{A}^{j}\mathbf{b}\in\mathcal{K}_{k}(\mathbf{A},\mathbf{b})$ for all $j<k$. Since $\mathbf{Q}\mathbf{Q}^{\mathsf{T}}$ is the 2-norm orthogonal projector onto $\mathcal{K}_{k}(\mathbf{A},\mathbf{b})$, then $\mathbf{Q}\mathbf{Q}^{\mathsf{T}}\mathbf{A}^{j}\mathbf{b}=\mathbf{A}^{j}\mathbf{b}$ for all $j<k$. Iteratively applying this fact,
$\mathbf{A}^{j}\mathbf{b}=\mathbf{Q}\mathbf{Q}^{\mathsf{T}}\mathbf{A}^{j}\mathbf{b}=\mathbf{Q}\mathbf{Q}^{\mathsf{T}}\mathbf{A}\mathbf{A}^{j-1}\mathbf{b}=\mathbf{Q}\mathbf{Q}^{\mathsf{T}}\mathbf{A}\mathbf{Q}\mathbf{Q}^{\mathsf{T}}\mathbf{A}^{j-1}\mathbf{b}=\mathbf{Q}\mathbf{Q}^{\mathsf{T}}\mathbf{A}\mathbf{Q}\cdots\mathbf{Q}^{\mathsf{T}}\mathbf{A}\mathbf{Q}\mathbf{Q}^{\mathsf{T}}\mathbf{b}.$
Using that $\mathbf{Q}^{\mathsf{T}}\mathbf{A}\mathbf{Q}=\mathbf{T}$, we find that $\mathbf{A}^{j}\mathbf{b}=\mathbf{Q}\mathbf{T}^{j}\mathbf{Q}^{\mathsf{T}}\mathbf{b}$, and thus, by linearity, $n(\mathbf{A})\mathbf{b}=\mathbf{Q}n(\mathbf{T})\mathbf{Q}^{\mathsf{T}}\mathbf{b}=\mathsf{lan}_{k}(n)$.
Lemma 2.4.
Let $\mathbf{A}_{j}$ be as in Eq. 9. Then, for any $k\geq 0$, $\mathsf{opt}_{k}(r_{j},\mathbf{A}_{j})=\mathbf{Q}(\mathbf{T}-z_{j}\mathbf{I})^{-1}\mathbf{Q}^{\mathsf{T}}r_{j-1}(\mathbf{A})\mathbf{b}$.
Proof 2.5.
The $\mathbf{A}_{j}$-norm projector onto $\mathcal{K}_{k}(\mathbf{A},\mathbf{b})$ is $\mathbf{Q}(\mathbf{Q}^{\mathsf{T}}\mathbf{A}_{j}\mathbf{Q})^{-1}\mathbf{Q}^{\mathsf{T}}\mathbf{A}_{j}$, so
$\mathsf{opt}_{k}(r_{j},\mathbf{A}_{j})=\mathbf{Q}(\mathbf{Q}^{\mathsf{T}}\mathbf{A}_{j}\mathbf{Q})^{-1}\mathbf{Q}^{\mathsf{T}}\mathbf{A}_{j}r_{j}(\mathbf{A}).$ (10)
Recall that $\mathbf{A}_{j}=\pm(\mathbf{A}-z_{j}\mathbf{I})$. Since $\mathbf{Q}^{\mathsf{T}}\mathbf{A}\mathbf{Q}=\mathbf{T}$ and $\mathbf{Q}^{\mathsf{T}}\mathbf{Q}=\mathbf{I}$, we have $\mathbf{Q}^{\mathsf{T}}\mathbf{A}_{j}\mathbf{Q}=\pm(\mathbf{Q}^{\mathsf{T}}(\mathbf{A}-z_{j}\mathbf{I})\mathbf{Q})=\pm(\mathbf{T}-z_{j}\mathbf{I})$. Noting also that $\mathbf{A}_{j}r_{j}(\mathbf{A})=\pm r_{j-1}(\mathbf{A})$ and plugging into Eq. 10 gives the result.
2.4 A telescoping sum for the error in terms of $\mathsf{opt}_{k}(r_{j},\mathbf{A}_{j})$
Our first main result is a decomposition for the Lanczos-FA error.
Lemma 2.6.
Suppose $k>\deg(n)$. Then,
$r(\mathbf{A})\mathbf{b}-\mathsf{lan}_{k}(r)=\sum_{j=1}^{q}\Bigg{[}\prod_{i=j+1}^{q}\mathbf{Q}(\mathbf{T}-z_{i}\mathbf{I})^{-1}\mathbf{Q}^{\mathsf{T}}\Bigg{]}\big{(}r_{j}(\mathbf{A})\mathbf{b}-\mathsf{opt}_{k}(r_{j},\mathbf{A}_{j})\big{)}.$
where we adopt the convention that $\prod_{i=q+1}^{q}\mathbf{B}_{i}=\mathbf{I}$ for any set of matrices $\{\mathbf{B}_{i}\}$.
Proof 2.7.
We can decompose $\mathsf{err}_{k}(r_{j})$ as
$\mathsf{err}_{k}(r_{j})$ $=r_{j}(\mathbf{A})\mathbf{b}-\mathbf{Q}r_{j}(\mathbf{T})\mathbf{Q}^{\mathsf{T}}\mathbf{b}$
$=\left[r_{j}(\mathbf{A})\mathbf{b}-\mathsf{opt}_{k}(r_{j},\mathbf{A}_{j})\right]+\left[\mathsf{opt}_{k}(r_{j},\mathbf{A}_{j})-\mathbf{Q}r_{j}(\mathbf{T})\mathbf{Q}^{\mathsf{T}}\mathbf{b}\right].$ (11)
Focusing on the second term and using Lemma 2.4,
$\mathsf{opt}_{k}(r_{j},\mathbf{A}_{j})-\mathbf{Q}r_{j}(\mathbf{T})\mathbf{Q}^{\mathsf{T}}\mathbf{b}$
$=\mathbf{Q}(\mathbf{T}-z_{j}\mathbf{I})^{-1}\mathbf{Q}^{\mathsf{T}}r_{j-1}(\mathbf{A})\mathbf{b}-\mathbf{Q}(\mathbf{T}-z_{j}\mathbf{I})^{-1}r_{j-1}(\mathbf{T})\mathbf{Q}^{\mathsf{T}}\mathbf{b}$
$=\mathbf{Q}(\mathbf{T}-z_{j}\mathbf{I})^{-1}\mathbf{Q}^{\mathsf{T}}\left[r_{j-1}(\mathbf{A})\mathbf{b}-\mathbf{Q}r_{j-1}(\mathbf{T})\mathbf{Q}^{\mathsf{T}}\mathbf{b}\right]$
$=\mathbf{Q}(\mathbf{T}-z_{j}\mathbf{I})^{-1}\mathbf{Q}^{\mathsf{T}}\left[\mathsf{err}_{k}(r_{j-1})\right].$
Substituting this into Eq. 11, we have
$\mathsf{err}_{k}(r_{j})=\left[r_{j}(\mathbf{A})\mathbf{b}-\mathsf{opt}_{k}(r_{j},\mathbf{A}_{j})\right]+\mathbf{Q}(\mathbf{T}-z_{j}\mathbf{I})^{-1}\mathbf{Q}^{\mathsf{T}}\mathsf{err}_{k}(r_{j-1}).$ (12)
Next, notice that $\mathsf{opt}_{k}(r_{1},\mathbf{A}_{1})$ is exactly the Lanczos approximation to $r_{1}(\mathbf{A})\mathbf{b}$. Indeed, since $\mathbf{Q}^{\mathsf{T}}\mathbf{Q}=\mathbf{I}$, Lemmas 2.2 and 2.4 imply
$\mathsf{lan}_{k}(r_{1})$ $=\mathbf{Q}(\mathbf{T}-z_{i}\mathbf{I})^{-1}n(\mathbf{T})\mathbf{Q}^{\mathsf{T}}\mathbf{b}$
$=\mathbf{Q}(\mathbf{T}-z_{i}\mathbf{I})^{-1}\mathbf{Q}^{\mathsf{T}}\mathbf{Q}n(\mathbf{T})\mathbf{Q}^{\mathsf{T}}\mathbf{b}$
$=\mathbf{Q}(\mathbf{T}-z_{1}\mathbf{I})^{-1}\mathbf{Q}^{\mathsf{T}}n(\mathbf{A})\mathbf{b}=\mathsf{opt}_{k}(r_{1},\mathbf{A}_{1}).$
Using this as a base case, we can repeatedly apply our decomposition of $\mathsf{err}_{k}(r_{j})$ in Eq. 12 to obtain
$\mathsf{err}_{k}(r)=\mathsf{err}_{k}(r_{q})=\sum_{j=1}^{q}\left[\prod_{i=j+1}^{q}\mathbf{Q}(\mathbf{T}-z_{i}\mathbf{I})^{-1}\mathbf{Q}^{\mathsf{T}}\right]\left(r_{j}(\mathbf{A})\mathbf{b}-\mathsf{opt}_{k}(r_{j},\mathbf{A}_{j})\right).$
2.5 Bounding each term in the telescoping sum and combining
The following lemma allows us to relate the optimality of the functions $r_{j}$ as measured in the $\mathbf{A}_{j}$-norm to that of $r_{q}=r$ as measured in the 2-norm.
Lemma 2.8.
For any $k>q-j$,
$\|r_{j}(\mathbf{A})\mathbf{b}-\mathsf{opt}_{k}(r_{j},\mathbf{A}_{j})\|$
$\leq\kappa(\mathbf{A}_{j})^{1/2}\cdot\|m_{j+1,q}(\mathbf{A})\|_{2}\min_{\mathrm{deg}(p)<k-(q-j)}\|r(\mathbf{A})\mathbf{b}-p(\mathbf{A})\mathbf{b}\|_{2}.$
Proof 2.9.
Recall that $\mathsf{opt}_{k}(r_{j},\mathbf{A}_{j})$ is the optimal approximation to $r_{j}(\mathbf{A})\mathbf{b}$ in the $\mathbf{A}_{j}$-norm, so can be related to the optimal approximation in the $2$-norm.
$\|r_{j}(\mathbf{A})\mathbf{b}-\mathsf{opt}_{k}(r_{j},\mathbf{A}_{j})\|_{2}$ $=\left\|\mathbf{A}_{j}^{-1/2}\mathbf{A}_{j}^{1/2}(r_{j}(\mathbf{A})\mathbf{b}-\mathsf{opt}_{k}(r_{j},\mathbf{A}_{j}))\right\|_{2}$
$\leq\|\mathbf{A}_{j}^{-1/2}\|_{2}\cdot\|r_{j}(\mathbf{A})\mathbf{b}-\mathsf{opt}_{k}(r_{j},\mathbf{A}_{j})\|_{\mathbf{A}_{j}}$
$=\|\mathbf{A}_{j}^{-1}\|_{2}^{1/2}\min_{\mathrm{deg}(p)<k}\|r_{j}(\mathbf{A})\mathbf{b}-p(\mathbf{A})\mathbf{b}\|_{\mathbf{A}_{j}}$
$\leq\|\mathbf{A}_{j}^{-1}\|_{2}^{1/2}\cdot\|\mathbf{A}_{j}\|_{2}^{1/2}\min_{\mathrm{deg}(p)<k}\|r_{j}(\mathbf{A})\mathbf{b}-p(\mathbf{A})\mathbf{b}\|_{2}$
$=\kappa(\mathbf{A}_{j})^{1/2}\min_{\mathrm{deg}(p)<k}\|r_{j}(\mathbf{A})\mathbf{b}-p(\mathbf{A})\mathbf{b}\|_{2}.$
We now relate the error of approximating $r_{j}(\mathbf{A})\mathbf{b}$ to that of approximating $r_{q}(\mathbf{A})\mathbf{b}=r(\mathbf{A})\mathbf{b}$:
$\min_{\mathrm{deg}(p)<k}\|r_{j}(\mathbf{A})\mathbf{b}-p(\mathbf{A})\mathbf{b}\|_{2}$
$=\min_{\mathrm{deg}(p)<k}\|m_{j+1,q}(\mathbf{A})r_{q}(\mathbf{A})\mathbf{b}-p(\mathbf{A})\mathbf{b}\|_{2}$
$\leq\min_{\mathrm{deg}(p)<k-(q-j)}\|m_{j+1,q}(\mathbf{A})r_{q}(\mathbf{A})\mathbf{b}-m_{j+1,q}(\mathbf{A})p(\mathbf{A})\mathbf{b}\|_{2}$
$\leq\|m_{j+1,q}(\mathbf{A})\|_{2}\min_{\mathrm{deg}(p)<k-(q-j)}\|r(\mathbf{A})\mathbf{b}-p(\mathbf{A})\mathbf{b}\|_{2}.$
Combining these two steps proves the lemma.
With the above results in place, we can prove Theorem 2.1. We will apply the triangle inequality to the telescoping sum for $\mathsf{err}_{k}(r)$ (Lemma 2.6) and bound each term by the error of the optimal approximation to $r(\mathbf{A})\mathbf{b}$ (Lemma 2.8).
Proof 2.10 (Proof of Theorem 2.1).
Focus on a single term in the sum in Lemma 2.6. First we will get rid of $\mathbf{Q}$ and $\mathbf{Q^{\mathsf{T}}}$. For $1\leq j<q$, using that $\mathbf{Q}^{\mathsf{T}}\mathbf{Q}=\mathbf{I}$ we have that
$\Bigg{\|}\prod_{i=j+1}^{q}\mathbf{Q}(\mathbf{T}-z_{i}\mathbf{I})^{-1}\mathbf{Q}^{\mathsf{T}}\Bigg{\|}_{2}$ $=\|\mathbf{Q}m_{j+1,q}(\mathbf{T})^{-1}\mathbf{Q}^{\mathsf{T}}\|_{2}\leq\|m_{j+1,q}(\mathbf{T})^{-1}\|_{2}.$ (13)
Also note that by the convention adopted in Lemma 2.6, for $j=q$, $\prod_{i=j+1}^{q}\mathbf{Q}(\mathbf{T}-z_{k}\mathbf{I})^{-1}\mathbf{Q}^{\mathsf{T}}=\mathbf{I}=m_{j+1,q}(\mathbf{T})$. Next we claim that
$\|m_{j+1,q}(\mathbf{T})^{-1}\|_{2}\leq\|m_{j+1,q}(\mathbf{A})^{-1}\|_{2}.$ (14)
To see why this is the case, note that because $\mathbf{T}=\mathbf{Q}^{\mathsf{T}}\mathbf{A}\mathbf{Q}$, the eigenvalues of $\mathbf{T}$ are contained in $\mathcal{I}=[\lambda_{\textup{min}}(\mathbf{A}),\lambda_{\textup{max}}(\mathbf{A})]$. By assumption, the roots of $m_{j+1,q}(x)$ are all real and lie outside of $\mathcal{I}$. Since there is at most one critical point between distinct roots, there can only be one critical point of $m_{j+1,q}(x)$ in $\mathcal{I}$. Thus, there can be no local minima of $|m_{j+1,q}(x)|$ in the interior of $\mathcal{I}$. Rather, the minimum of $|m_{j+1,q}(x)|$ over $\mathcal{I}$ must be attained at the boundary; i.e. at $x=\lambda_{\textup{min}}$ or $x=\lambda_{\textup{max}}$.
We now apply the triangle inequality to the telescoping sum of Lemma 2.6:
$\left\|\mathsf{err}_{k}(r)\right\|_{2}\leq\sum_{j=1}^{q}\Bigg{\|}\prod_{i=j+1}^{q}\mathbf{Q}(\mathbf{T}-z_{i}\mathbf{I})^{-1}\mathbf{Q}^{\mathsf{T}}\Bigg{\|}_{2}\cdot\big{\|}r_{j}(\mathbf{A})\mathbf{b}-\mathsf{opt}_{k}(r_{j},\mathbf{A}_{j})\big{\|}_{2}.$
Applying Eq. 13 and Eq. 14 we then have
$\left\|\mathsf{err}_{k}(r)\right\|_{2}\leq\sum_{j=1}^{q}\|m_{j+1,q}(\mathbf{A})^{-1}\|_{2}\cdot\big{\|}r_{j}(\mathbf{A})\mathbf{b}-\mathsf{opt}_{k}(r_{j},\mathbf{A}_{j})\big{\|}_{2}.$
Finally, using Lemma 2.8 we find
$\left\|\mathsf{err}_{k}(r)\right\|_{2}\leq\sum_{j=1}^{q}\kappa(\mathbf{A}_{j})^{1/2}\cdot\kappa(m_{j+1,q}(\mathbf{A}))\min_{\mathrm{deg}(p)<k-(q-j)}\|p(\mathbf{A})\mathbf{b}-r(\mathbf{A})\mathbf{b}\|_{2}.$ (15)
We can simplify by noting that
$\min_{\mathrm{deg}(p)<k-(q-j)}\|p(\mathbf{A})\mathbf{b}-r(\mathbf{A})\mathbf{b}\|_{2}\leq\min_{\mathrm{deg}(p)<k-(q-1)}\|p(\mathbf{A})\mathbf{b}-r(\mathbf{A})\mathbf{b}\|_{2},$
and we can combine all the condition number factors using
$\sum_{j=1}^{q}\kappa(\mathbf{A}_{j})^{1/2}\cdot\kappa(m_{j+1,q}(\mathbf{A}))\leq\sum_{j=1}^{q}\kappa(\mathbf{A}_{j})\prod_{i=j+1}^{q}\kappa(\mathbf{A}_{i})\leq q\prod_{i=1}^{q}\kappa(\mathbf{A}_{i}).$
The result follows by plugging into (15).
2.6 Comparison to Lanczos-OR
In [CGMM23] an algorithm called the Lanczos method for optimal rational function approximation (Lanczos-OR) was developed. For rational functions of the form Eq. 8, when run for $k>\deg(n)$ iterations, Lanczos-OR produces iterates
$\mathsf{lanOR}_{k}(r;\mathbf{A},\mathbf{b}):=\operatornamewithlimits{argmin}_{\mathbf{x}\in\mathcal{K}_{k-\lfloor q/2\rfloor}}\|r(\mathbf{A})\mathbf{b}-\mathbf{x}\|_{|r(\mathbf{A})|}.$
This implies a 2-norm instance optimality guarantee (Definition 1.1):
$\|r(\mathbf{A})\mathbf{b}-\mathsf{lanOR}_{k}(r;\mathbf{A},\mathbf{b})\|_{2}\leq\sqrt{\kappa(r(\mathbf{A}))}\min_{\deg(p)<k-\lfloor q/2\rfloor}\|r(\mathbf{A})\mathbf{b}-p(\mathbf{A})\mathbf{b}\|_{2}.$ (16)
This guarantee is similar to and often somewhat better than Theorem 2.1. For example, when $r(x)={1}/{m(x)}$, $\sqrt{\kappa(r(\mathbf{A}))}=\sqrt{\kappa(\mathbf{A}_{1})\cdot\kappa(\mathbf{A}_{2})\cdots\kappa(\mathbf{A}_{q})}$. In this case Eq. 16 improves on Theorem 2.1 by a $q$ factor and a square root. However, the bound is for a different algorithm; it does not extend to the ubiquitous Lanczos-FA method.
Moreover, despite this bound, Lanczos-OR usually performs worse than Lanczos-FA in practice if error is measured in the 2-norm. It is not hard to find examples where Eq. 16 is essentially sharp; see for instance Fig. 6. On the other hand, we have been unable to find any numerical examples where Theorem 2.1 is sharp, suggesting that it may be possible to prove a tighter bound for Lanczos-FA.
2.7 Implications for non-rational functions
In this section, we briefly discuss the implications of instance optimality guarantees for rational functions (like Theorem 2.1) to the problem of deriving guarantees for other (non-rational) functions. In particular, consider any function $f(x)$ that is uniformly well approximated by a low-degree rational function $r(x)$ on $\mathcal{I}=[\lambda_{\textup{min}},\lambda_{\textup{max}}]$, the field of values of $\mathbf{A}$. Furthermore, assume we have the following instance optimality guarantee for rational functions as in Definition 1.1:
$\|r(\mathbf{A})\mathbf{b}-\mathsf{lan}_{k}(r;\mathbf{A},\mathbf{b})\|_{2}\leq C_{r}\min_{\deg(p)<c_{r}k}\|r(\mathbf{A})\mathbf{b}-p(\mathbf{A})\mathbf{b}\|_{2}.$ (17)
We write the constants with subscripts to make it clear that they depend on the choice of approximant $r$. Theorem 2.1 is of this form, though as discussed above, its parameters may be loose. A natural way to approximate $f(\mathbf{A})\mathbf{b}$ would be to fix $r$, a rational approximation to $f$, and then apply Lanczos-FA to $r(\mathbf{A})\mathbf{b}$. However, the following analysis shows that Lanczos-FA applied to $f$ itself is nearly as good.
We can bound the error of Lanczos-FA on $f$ using triangle inequality as follows:
$\|f(\mathbf{A})\mathbf{b}-\mathsf{lan}_{k}(f;\mathbf{A},\mathbf{b})\|_{2}\leq\|f(\mathbf{A})\mathbf{b}-r(\mathbf{A})\mathbf{b}\|_{2}$ $+\|r(\mathbf{A})\mathbf{b}-\mathsf{lan}_{k}(r;\mathbf{A},\mathbf{b})\|_{2}$
$+\|\mathsf{lan}_{k}(r;\mathbf{A},\mathbf{b})-\mathsf{lan}_{k}(f;\mathbf{A},\mathbf{b})\|_{2}.$ (18)
The first and third terms of Subsection 2.7 are controlled by the maximum error of approximating $f$ with $r$ over $\mathcal{I}$. Specifically, if we let $\|r-f\|_{\mathcal{I}}:=\max_{x\in\mathcal{I}}|r(x)-f(x)|$,
$\|f(\mathbf{A})\mathbf{b}-r(\mathbf{A})\mathbf{b}\|_{2}\leq\|\mathbf{b}\|_{2}\cdot\|f(\mathbf{A})-r(\mathbf{A})\|_{2}\leq\|\mathbf{b}\|_{2}\cdot\|r-f\|_{\mathcal{I}}$
and similarly, using that $\Lambda(\mathbf{T})\subset\mathcal{I}$,
$\|\mathsf{lan}_{k}(r;\mathbf{A},\mathbf{b})-\mathsf{lan}_{k}(f;\mathbf{A},\mathbf{b})\|_{2}=\|\mathbf{Q}r(\mathbf{T})\mathbf{Q^{\mathsf{T}}b}-\mathbf{Q}f(\mathbf{T})\mathbf{Q^{\mathsf{T}}b}\|_{2}\\ \leq\|\mathbf{b}\|_{2}\cdot\|r-f\|_{\mathcal{I}}.$
The second term of Subsection 2.7 can controlled using Eq. 17 and a triangle inequality:
$\|r(\mathbf{A})\mathbf{b}-\mathsf{lan}_{k}(r;\mathbf{A},\mathbf{b})\|_{2}$ $\leq C_{r}\min_{\deg(p)<c_{r}k}\|r(\mathbf{A})\mathbf{b}-p(\mathbf{A})\mathbf{b}\|_{2}$
$\leq C_{r}\min_{\deg(p)<c_{r}k}\left(\|r(\mathbf{A})\mathbf{b}-f(\mathbf{A})\mathbf{b}\|_{2}+\|f(\mathbf{A})\mathbf{b}-p(\mathbf{A})\mathbf{b}\|_{2}\right)$
$\leq C_{r}\|\mathbf{b}\|_{2}\cdot\|r-f\|_{\mathcal{I}}+C_{r}\min_{\deg(p)<c_{r}k}\|f(\mathbf{A})b-p(\mathbf{A})\mathbf{b}\|_{2}.$
Combining, we obtain the bound
$\|f(\mathbf{A})\mathbf{b}-\mathsf{lan}_{k}(f;\mathbf{A},\mathbf{b})\|_{2}$
$\leq\min_{r}\Big{(}(C_{r}+2)\|\mathbf{b}\|_{2}\cdot\|r-f\|_{\mathcal{I}}+C_{r}\min_{\deg(p)<c_{r}k}\|f(\mathbf{A})\mathbf{b}-p(\mathbf{A})\mathbf{b}\|_{2}\Big{)}.$ (19)
Compare the form of this bound to that of Definitions 1.1 and 1.3. It is close to an instance near-optimality guarantee, expect for the first term, which requires $f$ to be uniformly well-approximated by a rational function $r$ on $\mathcal{I}=[\lambda_{\textup{min}},\lambda_{\textup{max}}]$. This is still much stronger than FOV optimality (Fact 1), which requires $f$ to be uniformly well-approximated by a polynomial to guarantee that Lanczos-FA provides a good approximation. There are many functions with lower degree rational approximations than polynomial approximations, even when we require the rational function $r$ to have poles only in $\mathbb{R}\setminus\mathcal{I}$ (as in our Theorem 2.1). For example, a uniform polynomial approximation to the square root on a strictly positive interval $[\lambda_{\textup{min}},\lambda_{\textup{max}}]$ requires degree $\Omega(\sqrt{\lambda_{\textup{max}}/\lambda_{\textup{min}}})$ [Tre19, Chapter 8]. On the other hand, a uniform rational approximation can be obtained with degree only $\Omega(\log(\lambda_{\textup{max}}/\lambda_{\textup{min}}))$ [HHT08, Tre19, Chapter 25]. Likewise, a uniform polynomial approximation to $\exp(-x)$ on the interval $[0,B]$ requires degree $\Omega(\sqrt{B})$ [AA22], but uniform rational approximations can be constructed with no dependence on $B$ at all [SSV75]. For such functions, we expect Eq. 19 to be stronger than Fact 1.
Notice also that, while we assume $f$ is well-approximated by a rational function, we are not required to actually construct the approximation. Indeed, since it holds for any $r$, instead of fixing a rational approximation of a certain degree, Eq. 19 automatically balances $\|r-f\|_{\mathcal{I}}$, which decreases as the degree grows, with $C_{r}$, which may increase as the degree grows. Currently, the instance optimality bound of Theorem 2.1 has a large $C_{r}$ that can grow exponentially with the degree of $r$. However, we conjecture that this dependence can be improved.
3 Spectrum optimality for $\mathbf{A}^{\pm 1/2}\mathbf{b}$
In this section, we prove spectrum optimality results for two especially important matrix functions, square root and inverse square root. We start by stating our result for the later.
Theorem 3.1.
Let $\Lambda$ be the spectrum of $\mathbf{A}$. Then for $k\geq 2$, the Lanczos-FA iterate satisfies the bound
$\|\mathbf{A}^{-1/2}\mathbf{b}-\mathsf{lan}_{k}(x^{-1/2};\mathbf{A},\mathbf{b})\|_{2}\leq\frac{3}{\sqrt{\pi k}}\kappa(\mathbf{A})\|\mathbf{b}\|_{2}\min_{\mathrm{deg}(p)<k/2}\left(\max_{x\in\Lambda}\left|\frac{1}{\sqrt{x}}-p(x)\right|\right).$
That is, Lanczos-FA applied to the inverse square root satisfies Definition 1.2 with
$C=\frac{3}{\sqrt{\pi k}}\kappa(\mathbf{A}),\qquad c=\frac{1}{2}.$
The proof of Theorem 3.1 relies on comparing the error of the $k$th Lanczos iterate for $x^{-1/2}$ to that of the Lanczos iterate for $x^{-1}$. First, applying a bound from [CGMM22], we use the Cauchy integral formula to upper bound the error of Lanczos-FA on $x^{\pm 1/2}$ by its error on $x^{-1}$ (Lemma 3.3). Second, as Equation Eq. 6 shows, Lanczos-FA is nearly instance optimal for the function $x^{-1}$; that is, it outputs $p(\mathbf{A})\mathbf{b}$ where $p$ is (nearly) the degree $k$ polynomial that best approximates $x^{-1}$. Third, the best degree $k$ polynomial approximation to $x^{-1}$ must have lower error than the best degree $k/2$ approximation to $x^{-1/2}$. This is because any degree $k/2$ approximation to $x^{-1/2}$ can be squared to yield a good degree $k$ polynomial approximation to $x^{-1}$. Combining these three steps upper bounds the error of the $k$th Lanczos-FA iterate for $x^{-1/2}$ by the error of the best degree $k/2$ polynomial approximation of $x^{-1/2}$.
Nearly the same argument can be used to prove spectrum optimality of Lanczos-FA for the function $\mathbf{A}^{-1/n}\mathbf{b}$ for any $n\in\mathbb{N}$ with $C=(2^{n}-1)\cdot\kappa(\mathbf{A})/\sqrt{\pi k}$. Furthermore, recall that using Lemma 1.4, we can convert Theorem 3.1 into a near instance optimality guarantee at the price of strong dependence of $C$ on $\mathbf{b}$.
We next state our optimality result for the matrix square root.
Theorem 3.2.
Let $\Lambda$ be the spectrum of $\mathbf{A}$. Then for $k\geq 2$, the Lanczos-FA iterate satisfies the bound
$\|\mathbf{A}^{1/2}\mathbf{b}-\mathsf{lan}_{k}(x^{1/2};\mathbf{A},\mathbf{b})\|\leq\frac{3\kappa(\mathbf{A})^{2}}{k^{3/2}}\|\mathbf{b}\|_{2}\min_{\mathrm{deg}(p)<k/2+1}\left(\max_{x\in\Lambda\cup\{0\}}\left|\sqrt{x}-p(x)\right|\right).$
This bound resembles Definition 1.2 with
$C=\frac{3\kappa(\mathbf{A})^{2}}{k^{3/2}},\qquad c=\frac{1}{2}.$
However, it is slightly weaker in that the maximization is taken over $\Lambda\cup\{0\}$ instead of only $\Lambda$.
The proof of Theorem 3.2 is nearly the same as that of Theorem 3.1. Ideally, if $p$ is a polynomial approximation to $x^{1/2}$, we would like to claim that $(p(x)/x)^{2}$ yields a good polynomial approximation to $x^{-1}$. However, since this function is not necessarily a polynomial, we must instead use use $\left(\frac{p(x)-p(0)}{x}\right)^{2}$, which introduces the need to include $\{0\}$ in the maximization on the right-hand side.
Figure 3: The bounds of Theorems 3.1 and 3.2 capture the convergence behavior of Lanczos-FA for $\mathbf{A}^{\pm 1/2}\mathbf{b}$. In particular, they can be tighter than the FOV optimality bound of Fact 1. The predicted rate of convergence is about half that observed for Lanczos-FA in this example; this is due to the parameter $c=1/2$ in both bounds.
3.1 Proofs
We begin by quoting bounds from [CGMM22].
Lemma 3.3.
For all $k\geq 1$, the Lanczos-FA iterate satisfies the bounds
  1. 1.
    $\|\mathbf{A}^{1/2}\mathbf{b}-\mathsf{lan}_{k}(x^{1/2};\mathbf{A},\mathbf{b})\|_{2}\leq\frac{\lambda_{\textup{max}}^{3/2}}{2k^{3/2}}\cdot\|\mathbf{A}^{-1}\mathbf{b}-\mathsf{lan}_{k}(1/x;\mathbf{A},\mathbf{b})\|_{2}.$
  2. 2.
    $\|\mathbf{A}^{-1/2}\mathbf{b}-\mathsf{lan}_{k}(x^{-1/2};\mathbf{A},\mathbf{b})\|_{2}\leq\sqrt{\frac{\lambda_{\textup{max}}}{\pi k}}\cdot\|\mathbf{A}^{-1}\mathbf{b}-\mathsf{lan}_{k}(1/x;\mathbf{A},\mathbf{b})\|_{2}.$
Proof 3.4.
For Item 1, see Example 4.1 in [CGMM22]. For part Item 2, we use the same proof again but substituting the inverse square root for the square root. For both parts, we have simplified the resulting bounds using the fact that for $k\geq 1$,
$\frac{\Gamma(k-1/2)}{\Gamma(k+1)}\leq\frac{\sqrt{\pi}}{k^{3/2}}\qquad\qquad\frac{\Gamma(k+1/2)}{\Gamma(k+1)}\leq\frac{1}{\sqrt{k}}.$
Next, we bound the error of Lanczos-FA on linear systems by the error of the optimal polynomial approximation to $x^{-1/2}$:
Lemma 3.5.
For all $k\geq 1$, the Lanczos-FA iterate satisfies the bounds
$\|\mathbf{A}^{-1}\mathbf{b}-\mathsf{lan}_{k}(1/x;\mathbf{A},\mathbf{b})\|_{2}\leq\frac{3}{\sqrt{\lambda_{\textup{min}}}}\sqrt{\kappa(\mathbf{A})}\|\mathbf{b}\|_{2}\min_{\mathrm{deg}(p)<k/2}\left(\max_{x\in\Lambda}\left|\frac{1}{\sqrt{x}}-p(x)\right|\right).$
Proof 3.6.
Let
$p^{*}(x)=\operatornamewithlimits{argmin}_{\mathrm{deg}(p)<k/2}\max_{x\in\Lambda}\left|\frac{1}{\sqrt{x}}-p(x)\right|$
and note that $p(x)^{2}$ is a polynomial of degree less than $k$. Therefore,
$\min_{\deg(p)<k}\|\mathbf{A}^{-1}\mathbf{b}-p(\mathbf{A})\mathbf{b}\|_{2}\leq\|\mathbf{b}\|_{2}\min_{\deg p<k}\max_{x\in\Lambda}\left|\frac{1}{x}-p(x)\right|\leq\|\mathbf{b}\|_{2}\max_{x\in\Lambda}\left|\frac{1}{x}-p^{*}(x)^{2}\right|.$
Thus, since $|1/x-p^{*}(x)^{2}|=|1/\sqrt{x}+p^{*}(x)|\cdot|1/\sqrt{x}-p^{*}(x)|$, we can plug the above into Eq. 6 to obtain a bound
$\|\mathbf{A}^{-1}\mathbf{b}-$ $\mathsf{lan}_{k}(1/x)\|_{2}\leq\sqrt{\kappa(\mathbf{A})}\|\mathbf{b}\|_{2}\left(\max_{x\in\Lambda}\left|\frac{1}{\sqrt{x}}+p^{*}(x)\right|\right)\left(\max_{x\in\Lambda}\left|\frac{1}{\sqrt{x}}-p^{*}(x)\right|\right).$
Now use can use the optimality of $p^{*}(x)$in approximating $1/\sqrt{x}$ on $\Lambda$ to bound:
$\max_{x\in\Lambda}\left|\frac{1}{\sqrt{x}}+p^{*}(x)\right|$ $=\max_{x\in\Lambda}\left|\frac{2}{\sqrt{x}}+p^{*}(x)-\frac{1}{\sqrt{x}}\right|$
$\leq\max_{x\in\Lambda}\left|\frac{2}{\sqrt{x}}\right|+\max_{x\in\Lambda}\left|p^{*}(x)-\frac{1}{\sqrt{x}}\right|$
$\leq\frac{2}{\sqrt{\lambda_{\textup{min}}}}+\max_{x\in\Lambda}\left|0(x)-\frac{1}{\sqrt{x}}\right|\leq\frac{3}{\sqrt{\lambda_{\textup{min}}}}.$
Substituting this inequality above proves the lemma.
The main proofs now follow directly.
Proof 3.7 (Proof of Theorem 3.1).
Substitute Lemma 3.5 into Lemma 3.3, part (b).
Proof 3.8 (Proof of Theorem 3.2).
Let
$p^{*}(x)=\operatornamewithlimits{argmin}_{\mathrm{deg}(p)<k/2+1}\max_{x\in\Lambda\cup\{0\}}\left|\sqrt{x}-p(x)\right|,\qquad\tilde{p}(x)=\frac{p^{*}(x)-p^{*}(0)}{x}$
and note that $\tilde{p}$ is a polynomial of degree less than $k/2$. Then
$\min_{\mathrm{deg}(p)<k/2}\left(\max_{x\in\Lambda}\left|\frac{1}{\sqrt{x}}-p(x)\right|\right)$ $\leq\max_{x\in\Lambda}\left|\frac{1}{\sqrt{x}}-\tilde{p}(x)\right|$
$=\max_{x\in\Lambda}\left(\left|1/x\right|\cdot\left|\sqrt{x}-x\tilde{p}(x)\right|\right)$
$\leq\frac{1}{\lambda_{\textup{min}}}\max_{x\in\Lambda}\left|\sqrt{x}-p^{*}(x)+p^{*}(0)\right|$
$\leq\frac{1}{\lambda_{\textup{min}}}\max_{x\in\Lambda}\left(\left|\sqrt{x}-p^{*}(x)\right|+\left|\sqrt{0}-p^{*}(0)\right|\right)$
$\leq\frac{2}{\lambda_{\textup{min}}}\max_{x\in\Lambda\cup\{0\}}\left|\sqrt{x}-p^{*}(x)\right|.$
Combining Lemma 3.3 part (a), Lemma 3.5, and the above proves the theorem.
4 Experiments
We implemented Lanczos-FA in high precision arithmetic using the flamp library, which is based on mpmath [${}^{+}$13]. [^5] Code for all our experiments is available at https://github.com/NoahAmsel/lanczos-optimality/tree/near_optimality_paper. In Fig. 1, we compare its performance to the instance optimal, spectrum optimal, and FOV optimal approximations (the right hand sides of Definitions 1.1, 1.2, and 1.3 with $C=c=1$) for various matrix functions and spectra. We use three test matrices $\mathbf{A}\in\mathbb{R}^{100\times 100}$ which all have condition number $100$. The first has a uniformly-spaced spectrum, the second has a geometrically-spaced spectrum, and the third has eigenvalues that are all uniformly-spaced in a narrow interval except for ten very small eigenvalues. As suggested by Lemma 1.4, if $\mathbf{b}$ is chosen randomly, the spectrum optimal error nearly coincides with the instance optimal error; to demonstrate that they can differ significantly, we choose $\mathbf{b}$ to be a linear combination of $\mathbf{A}$’s eigenvectors, with more weight given to eigenvectors with larger eigenvalues. We construct the FOV and spectrum optimal approximations using the Remez algorithm, and the instance optimal approximation using least squares regression. In these experiments, as in almost all cases we tried, Lanczos-FA performs nearly as well as the instance optimal approximation. As shown in the bottom panel in Fig. 1, its error is never more than $2.5$x the optimal error. While FOV optimality is a standard way of understanding Lanczos-FA’s convergence (Fact 1), it fails to capture the true convergence behavior in all of these cases.
To understand Theorem 2.1, we compare the bound to the true convergence curve of Lanczos-FA for various rational functions in Fig. 2. We also plot the FOV optimality bound of Fact 1. We use the same matrices and $\mathbf{b}$ vectors as in Fig. 1; results are similar if $\mathbf{b}$ is instead chosen as a uniform linear combination of $\mathbf{A}$’s eigenvectors. We choose rational functions to match the functions used for Fig. 1. We construct a degree 5 rational approximation to $\exp(-x/10)$ following [SSV75]. We construct a degree 10 approximation to $\log(x)$ using the BRASIL algorithm [Hof21] and verify that it has real poles outside the interval of the eigenvalues. Despite Theorem 2.1’s exponential dependence on the degree of the rational function being applied, Fig. 2 shows that it matches the shape of the convergence curve well and is tighter than Fact 1 when the number of iterations is large. That said, in all cases plotted, Lanczos-FA always returns an approximation within a factor $4$ of optimal, suggesting that the leading coefficient in Theorem 2.1 is not nearly tight.
To understand Theorems 3.1 and 3.2, we repeat the experiment using the inverse square root and square root functions. We reuse the second and third $\mathbf{A}$ matrix and $\mathbf{b}$ vectors from the previous experiment. As Fig. 3 shows, our bounds are tighter than Fact 1. They closely resemble the true convergence of Lanczos-FA, but stretched horizontally by a factor of 2. This is because the degree of the polynomial minimization in both bounds is $k/2$, while in these examples Lanczos-FA performs nearly instance optimally (that is, as well as the best degree $k$ polynomial). Still, our bounds capture the correct qualitative behavior of the algorithm.
4.1 Dependence on the rational function degree
Figure 4: The maximum observed ratio between the error of Lanczos-FA and the optimal error over choice of right hand side $\mathbf{b}$ when approximating $\mathbf{A}^{-q}$ for matrices with varying condition number $\kappa$. Each point shows the optimality ratio for a different pair of $\kappa$ and $q$. Points with the same color correspond to the same value of $\kappa$. On the left, the dotted line plots $\sqrt{q\kappa}$ for the maximum $\kappa$ considered ($10^{6}$). On the right, the dotted line plots $\sqrt{q\kappa}$ for the maximum $q$ considered ($2^{6}$). Overall, the optimality ratio appears to scale at least as $\Omega(\sqrt{q\kappa})$. While it is possible that a more refined method for finding hard problem instances could result in a strong lower bound, we conjecture that the $q\kappa^{q}$ ratio proven in Theorem 2.1 for this problem is far from tight.
Theorem 2.1 upper bounds the optimality ratio by $C=q\cdot\kappa(\mathbf{A}_{1})\cdot\kappa(\mathbf{A}_{2})\cdots\kappa(\mathbf{A}_{q})\leq q\gamma^{q}$, where $\gamma$ is as defined in Subsection 1.3. We conjecture that this bound can be tightened. However, the following experiment provides evidence that it is not possible to entirely eliminate the dependence on the rational function’s denominator degree $q$. In particular, for parameters $(\kappa,q)$, consider approximating $\mathbf{A}^{-q}\mathbf{b}$ where $\mathbf{A}$ has spectrum $\lambda_{1}=1$ and $\lambda_{2},\ldots,\lambda_{100}$ evenly spaced between $0.999995\cdot\kappa$ and $\kappa$. In this case, $\kappa(\mathbf{A}_{1})=\cdots=\kappa(\mathbf{A}_{q})=\kappa(\mathbf{A})=\kappa$. We generate a grid of problems by picking different combinations of $(\kappa,q)$ and tuning $\mathbf{b}$ in a limited way to maximize the maximum ratio between the error of Lanczos-FA and the optimal Krylov error over all iterations. In particular, we took $\mathbf{b}$ to be an all ones vector, except we varied its first entry, using grid search to maximize the optimality ratio. The results, plotted in Fig. 4, suggest that the optimality ratio grows at least as $\Omega(\sqrt{q\cdot\kappa(\mathbf{A})})$.
4.2 Rational functions with poles between eigenvalues
In Theorem 2.1, we assume that the poles of the rational function $r$ lie outside the interval $\mathcal{I}=[\lambda_{\textup{min}},\lambda_{\textup{max}}]$ containing $\mathbf{A}$’s eigenvalues. If the poles of $r$ live in $\mathcal{I}$, then the Lanczos-FA iterate can be arbitrarily far from the optimal iterate. Indeed, an eigenvalue of $\mathbf{T}$ might be very close to a pole, causing $f(\mathbf{T})$ to be poorly behaved. This behavior is easily observed e.g. for $f(\mathbf{A})=\mathbf{A}^{-1}$ and rules out a direct analog of Theorem 2.1 when $r(x)$ has poles in $\mathcal{I}$.
However, as discussed in Subsection 1.2, the “overall” convergence of Lanczos-FA still seems to closely follow the optimal approximation for many matrix functions. Specifically, even if the Lanczos-FA iterate deviates significantly from optimal on some iterations, there are other iterations where it is very close to optimal. We illustrate this phenomena in Fig. 5 for two rational functions with poles in $\mathcal{I}=[\lambda_{\textup{min}},\lambda_{\textup{max}}]$, as well as the function $\operatorname{sign}(x)$, which has a discontinuity in $\mathcal{I}$ and would typically be approximated by a rational function with conjugate pairs of imaginary poles on the imaginary axis which become increasingly close to the the interval as the rational function degree increases. [^6] We remark that [CGMM23] shows that applying their Lanczos-OR method to each term in the partial fraction decomposition of the approximating rational function yields an approximation to the sign function that seems to avoid the oscillations of Lanczos-FA.
Figure 5: Convergence of Lanczos-FA for an indefinite matrix. The spectrum of $\mathbf{A}$ is symmetric about 00 and geometrically spaced between $-100$ and $-1$ and between $1$ and $100$. When the rational function has poles that lie in the interval of $\mathbf{A}$’s eigenvalues or that are imaginary, the optimality ratio can be very large for some iterations. Similar behavior is seen for functions like $\operatorname{sign}(\mathbf{A})$ that have a discontinuity in the interval of $\mathbf{A}$’s eigenvalues. However, the “overall” convergence of Lanczos-FA still appears to closely track the instant optimal solution.
Developing a better theoretical understanding of this observed “overall” near-optimality for functions where Lanczos-FA behaves poorly on the worst-case iterations is an interesting direction for future study. In Subsection A.3 we make an initial step by providing an explanation of the phenomenon for the function $\mathbf{A}^{-1}$ when $\mathbf{A}$ is indefinite, drawing on previous work on the convergence of MINRES and CG for indefinite systems [CG96, GDK99].
4.3 Comparison with Lanczos-OR
As discussed in Subsection 2.6, the Lanczos-OR method is an alternative method for rational matrix function approximation for which a tighter guarantee than Theorem 2.1 is currently known (the iterates are optimal in a certain norm). However, Lanczos-FA usually outperforms Lanczos-OR when measured in the 2-norm. For example, as discussed in Subsection 4.1, when approximating $\mathbf{A}^{-q}\mathbf{b}$, the worst optimality ratio of Lanczos-FA that we were able to observe was $O(\sqrt{q\cdot\kappa(\mathbf{A})})$, much lower than Theorem 2.1 would suggest). Repeating the same experiment with Lanczos-OR shows that the method’s optimality ratio appears to grow as $\Omega(\kappa(\mathbf{A})^{q/2})$ (see Fig. 6).
Figure 6: The maximum observed ratio between the error of Lanczos-OR and the optimal error over choice of right hand side $\mathbf{b}$ when approximating $\mathbf{A}^{-q}$ for matrices with varying condition number $\kappa$. Each point shows the optimality ratio for a different pair of $\kappa$ and $q$. Points with the same color correspond to the same value of $\kappa$. On the left, the dotted line plots $g(q)=\kappa^{q/2}$ for the maximum $\kappa$ considered ($10^{6}$). On the right, the dotted line plots $g(\kappa)=\kappa^{q/2}$ for the maximum $q$ considered ($642^{6}=$ 64). Overall, the optimality ratio appears to grow as $\Omega(\kappa^{q/2})$. Contrast with Fig. 4.
4.4 Non-rational functions
As noted in Subsection 2.7, an instance optimality guarantee for rational functions also implies that Lanczos-FA performs well on functions that are well-approximated by rationals. As an example, we consider the function is $f(x)=x^{-0.4}$, for which a rational approximation in any degree can be found using the BRASIL algorithm [Hof21]. Figure 7 shows how applying Lanczos-FA to these rational approximations compares to applying Lanczos-FA directly to the $f$ itself to approximate $\mathbf{A}^{-0.4}\mathbf{b}$. When the number of iterations is small, both methods perform nearly optimally, as the accuracy is limited more by the small size of the Krylov subspace than by the difference between $f$ and the rational approximant (that is, the second term in Eq. 19 dominates the first term). As the number of iterations grows, the error due to approximating $f$ in the Krylov subspace continues to decrease while the error of uniformly approximating $f$ by some fixed rational function remains fixed (that is, the first term in Eq. 19 dominates the first term); however, increasing the degree of the rational approximant decreases this source of error. This shows that understanding the convergence of Lanczos-FA for the entire family of rational approximations goes a long way toward explaining its convergence for non-rational functions. In the limit, Lanczos-FA applied to $f$ itself appears to automatically inherit the instance optimality of a suitably high-degree rational approximation.
Figure 7: Applying Lanczos-FA to the function $\mathbf{A}^{-0.4}$ and rational approximations of various degrees found using the BRASIL algorithm [Hof21]. In this experiment, the spectrum of $\mathbf{A}$ contains two clusters: 10 eigenvalues uniformly spaced near 1, and 90 eigenvalues uniformly spaced near 100. As predicted by the bound in Subsection 2.7, convergence of Lanczos-FA for this function appears to closely track that of a high degree rational approximant.
5 Outlook
This paper provides instance optimality guarantees for Lanczos-FA applied to a broad class of rational functions, as well as weaker spectrum optimality guarantees for $\mathbf{A}^{1/2}$ and $\mathbf{A}^{-1/2}$. While this is, to the best of our knowledge, the first time Lanczos-FA has been shown to be nearly instance optimal for a function besides $\mathbf{A}^{-1}$, our analysis is only a small step towards truly understanding the remarkable behavior of Lanczos-FA. In the remainder of this section, we provide an overview of some research directions which we believe are worthy of further study.
Extension to other function classes
Empirically, Lanczos-FA seems to attain near-spectrum or near-instance optimality for a wide variety of functions beyond those considered in this paper. For example, rational functions with conjugate pairs of complex poles whose real parts lie outside the interval of the eigenvalues seem to behave similarly to those with strictly real poles, which we study in Theorem 2.1. It would be interesting to extend our results to such settings. Moreover, improving on the results of [DGK98], we conjecture that near spectrum optimality holds for the exponential function with a $C$ that depends only on $\lambda_{\textup{min}}$ and $\lambda_{\textup{max}}$.
Construction of hard instances / refined upper bounds
Theorem 2.1 has an exponential dependence on the degree of the rational function’s denominator $q$ which limits the practicality of our bounds. It is unclear if/when this dependence can be improved. In fact, even for the task of approximating $\mathbf{A}^{-q}\mathbf{b}$ for positive definite $\mathbf{A}$, and despite having studied this particular problem for over two years, the authors still have differing beliefs as to whether the optimality ratio of Lanczos-FA should have an exponential dependence on $q$. The experiment in Subsection 4.1 provides strong evidence that some dependence on $q$ is necessary, but the hardest examples we have depend on $\sqrt{q}$, instead of the current bound of $\kappa(\mathbf{A})^{q}$ guaranteed by Theorem 2.1. It would be great to either tighten the bound of Theorem 2.1, or find hard instances which show that no stronger bound can be proven for Lanczos-FA.
We note that, to study the above problem, it might be fruitful to consider other norms besides the $2$-norm that appears in Definition 1.1 and Theorem 2.1. For instance, when measured in the $\mathbf{A}$-norm, Lanczos-FA on $\mathbf{A}^{-q}$ does not exhibit the same $\sqrt{\kappa}$ gap from optimal observed in Subsection 4.1. This suggests that, for this class of functions, the $\mathbf{A}$-norm is more natural. As in work on the Lanczos-OR method [CGMM23], it may be easier to prove strong near optimality bounds for different norms, perhaps depending on $r$.
As discussed in Subsection 4.2, it would also be interesting to provide bounds for general rational functions (possibly with poles inside the interval $[\lambda_{\textup{max}},\lambda_{\textup{min}}]$) akin to our bounds in Subsection A.3 for $f(x)=1/x$ when $\mathbf{A}$ is indefinite. The behavior of Lanczos-FA on such problems is often extremely interesting: at some iterations the error is huge, but the minimum error observed up to a given iteration typically seems to be nearly the error of the optimal Krylov subspace method.
Connection to approximation
Let $\mathbf{u}_{1},\ldots,\mathbf{u}_{n}$ be the orthogonal eigenvectors of a symmetric matrix $\mathbf{A}$. Define a weight function
$\omega(x):=\sum_{i=1}^{d}\delta_{\lambda_{i}}(x)|\mathbf{u}_{i}^{\mathsf{T}}\mathbf{b}|^{2},$
where $\delta_{c}$ is the Dirac distribution centered at $c$. Such a function is often referred to as the eigenvector spectral distribution (VESD) for $(\mathbf{A},\mathbf{b})$. It is well-known that the eigenvalues of $\mathbf{T}$ are exactly the zeros of the degree $k+1$ orthogonal polynomial of $\omega$. This implies that $\mathsf{lan}_{k}(f;\mathbf{A},\mathbf{b})=\tilde{p}(\mathbf{A})\mathbf{b}$, where $\tilde{p}$ is the degree $k-1$ polynomial interpolant to $f$ at the eigenvalues of $\mathbf{T}$. Thus, understanding $\|f(\mathbf{A})\mathbf{b}-\mathsf{lan}_{k}(f;\mathbf{A},\mathbf{b})\|_{2}$ is equivalent to understanding the $\omega$ weighted $L^{2}$ error $\int|f(x)-\tilde{p}(x)|^{2}\omega(x)\mathrm{d}{x}$ of the polynomial interpolant $\tilde{p}$ to $f$ at the zeros of the degree $k$ orthogonal polynomial of $\omega$. This general perspective may allow ideas from approximation theory to be applied to this problem.
Finite precision arithmetic
The present analysis deals only with the mathematical behavior of the Lanczos algorithm; that is, the behavior when the algorithm is run in exact arithmetic. In practice, the implementation of the Lanczos algorithm is very important; in particular, the most efficient implementations, which do not explicitly orthogonalize the basis vectors, are highly susceptible to the impacts of finite precision arithmetic and output $\mathbf{Q}$ which are far from orthogonal [MS06, CLS22]. While this instability can be more or less mitigated with (more expensive) explicit orthogonalization schemes, there is a lot of theoretical work showing that Lanczos and Lanczos-FA still work well without reorthogonalization [Pai71, Pai76, Pai80, Gre89, DK92, DGK98]. For example, [DK92, MMS18] show that the FOV bound of Fact 1 still holds up to a close approximation in finite precision arithmetic for any bounded matrix function. It would be interesting to study whether stronger near-optimality guarantees like those proven in Theorem 2.1 are also robust to finite precision. For $f(x)=1/x$, this problem has already been studied in [Gre89] and related work.
Appendix A Deferred proofs
In this section we provide proofs for two facts mentioned in the introduction. While both facts are easy corollaries of past work, we believe that are of independent interest.
A.1 Instance, spectrum, and FOV optimality can coincide
Theorem A.1.
  1. 1.
    Given any $d$, discrete set $\Lambda$ of at most $d$ real points, $f$, and $k<d$, there exists a symmetric matrix $\mathbf{A}\in\mathbb{R}^{d\times d}$ and vector $\mathbf{b}\in\mathbb{R}^{d}$ such that the spectrum of $\mathbf{A}$ is $\Lambda$ and
    $\min_{\deg(p)<k}\|f(\mathbf{A})\mathbf{b}-p(\mathbf{A})\mathbf{b}\|_{2}/{\|\mathbf{b}\|_{2}}=\min_{\deg(p)<k}\max_{x\in\Lambda}|f(x)-p(x)|.$
  2. 2.
    Given any $d$, $\lambda_{\textup{min}}$, $\lambda_{\textup{max}}$, $f$ continuous on $[\lambda_{\textup{min}},\lambda_{\textup{max}}]$, and $k<d$, there exists a symmetric matrix $\mathbf{A}\in\mathbb{R}^{d\times d}$ and vector $\mathbf{b}\in\mathbb{R}^{d}$ such that the smallest eigenvalue of $\mathbf{A}$ is $\lambda_{\textup{min}}$, the largest eigenvalue of $\mathbf{A}$ is $\lambda_{\textup{max}}$, and
    $\min_{\deg(p)<k}\|f(\mathbf{A})\mathbf{b}-p(\mathbf{A})\mathbf{b}\|_{2}/{\|\mathbf{b}\|_{2}}=\min_{\deg(p)<k}\max_{x\in\mathcal{I}}|f(x)-p(x)|.$
Our proof follows the approach of [Gre79, Theorem 1] closely.
Proof A.2.
For Item 1, note that if $|\Lambda|\leq k$, then both sides of the expression are zero. Thus, it suffices to consider the case $|\Lambda|\geq k+1$. On a discrete set $\Lambda$ with at least $k+1$ distinct points, $f$ has a best (on $\Lambda$) polynomial approximation $p^{*}$ of degree $k-1$ that equioscillates $f$ at $k+1$ points [Riv81, Theorem 1.11]. That is, there exist values $\lambda_{1}<\cdots<\lambda_{k+1}$ contained in $\Lambda$ such that
$f(\lambda_{i})-p^{*}(\lambda_{i})=(-1)^{i-1}\epsilon,\qquad i=1,\ldots,k+1,$
where
$\epsilon:=\max_{x\in\Lambda}|f(x)-p^{*}(x)|.$
Similarly, for Item 2 if $f$ is continuous on $\mathcal{I}=[\lambda_{\textup{min}},\lambda_{\textup{max}}]$, then $f$ has a unique best polynomial approximation $p^{*}$ of degree $k-1$ which equioscillates at $k+1$ points $\lambda_{1}<\cdots<\lambda_{k+1}$ in $\mathcal{I}$ (see for instance [Riv81, Tre19]), and we analogously define $\epsilon$ as
$\epsilon:=\max_{x\in\mathcal{I}}|f(x)-p^{*}(x)|.$
Set
[…]
Then,
$\min_{\deg(p)<k}\|f(\mathbf{A})\mathbf{b}-p(\mathbf{A})\mathbf{b}\|_{2}^{2}=\min_{\alpha_{0},\ldots,\alpha_{k-1}}\sum_{i=1}^{k+1}b_{i}^{2}\big{(}f(\lambda_{i})-p(\lambda_{i})\big{)}^{2},\qquad p(\lambda):=\sum_{j=0}^{k-1}\alpha_{j}\lambda^{j}.$
We would like to find $\mathbf{b}$ so that $f(\lambda_{i})-p(\lambda_{i})=(-1)^{i-1}\epsilon$; i.e. so that $p=p^{*}$. When $\mathbf{b}$ is fixed, for each $j=0,1,\ldots,k-1$ the solution to this least squares problem must satisfy
$10=\frac{\mathrm{d}}{\mathrm{d}\alpha_{j}}\sum_{i=1}^{k+1}b_{i}^{2}\big{(}f(\lambda_{i})-p(\lambda_{i})\big{)}^{2}=-2\sum_{i=1}^{k+1}b_{i}^{2}\left(f(\lambda_{i})-p(\lambda_{i})\right)\lambda_{i}^{j},\qquad j=0,1,\ldots,k-$ 1 .
When $p=p^{*}$, this gives the conditions
$10=-2\sum_{i=1}^{k+1}b_{i}^{2}(-1)^{i-1}\epsilon\lambda_{i}^{j},\qquad j=0,1,\ldots,k-$ 1 .
Without loss of generality, we can assume $\|\mathbf{b}\|_{2}=1$ so that $b_{1}^{2}+\cdots+b_{k+1}^{2}=1$. Thus, we obtain a linear system
$\begin{bmatrix}1&1&\cdots&1\\ -2\epsilon\lambda_{1}^{0}&2\epsilon\lambda_{2}^{0}&\cdots&2(-1)^{k+1}\lambda_{k+1}^{0}\epsilon\\ \vdots&\vdots&&\vdots\\ -2\epsilon\lambda_{1}^{k-1}&2\epsilon\lambda_{2}^{k-1}&\cdots&2(-1)^{k+1}\epsilon\lambda_{k+1}^{k-1}\\ \end{bmatrix}\begin{bmatrix}b_{1}^{2}\\ b_{2}^{2}\\ \vdots\\ b_{k+1}^{2}\end{bmatrix}=\begin{bmatrix}1\\ 0\\ \vdots\\ 0\end{bmatrix}.$
We can rewrite this system as
$\begin{bmatrix}1&-1&\cdots&(-1)^{k}\\ \lambda_{1}^{0}&\lambda_{2}^{0}&\cdots&\lambda_{k+1}^{0}\\ \vdots&\vdots&&\vdots\\ \lambda_{1}^{k-1}&\lambda_{2}^{k-1}&\cdots&\lambda_{k+1}^{k-1}\\ \end{bmatrix}\begin{bmatrix}b_{1}^{2}\\ -b_{2}^{2}\\ \vdots\\ (-1)^{k}b_{k+1}^{2}\end{bmatrix}=\begin{bmatrix}1\\ 0\\ \vdots\\ 0\end{bmatrix}.$
This system can be solved analytically via Cramer’s rule and has solution
$b_{\ell}^{2}=\Bigg{(}\prod_{\begin{subarray}{c}i=1\\ i\neq\ell\end{subarray}}^{k+1}\prod_{\begin{subarray}{c}j=i+1\\ j\neq\ell\end{subarray}}^{k+1}(\lambda_{j}-\lambda_{i})\Bigg{)}\Bigg{/}\Bigg{(}\sum_{m=1}^{k+1}\prod_{\begin{subarray}{c}i=1\\ i\neq m\end{subarray}}^{k+1}\prod_{\begin{subarray}{c}j=i+1\\ j\neq m\end{subarray}}^{k+1}(\lambda_{j}-\lambda_{i})\Bigg{)}.$
Clearly $(\lambda_{j}-\lambda_{i})$ is positive for $j\in\{i+1,\ldots k+1\}$. Thus the numerator and denominator above are positive, so the entries of $\mathbf{b}$ are well-defined and real. This implies that, for these choices of $\mathbf{A}$ and $\mathbf{b}$,
$\min_{\deg(p)<k}\|f(\mathbf{A})\mathbf{b}-p(\mathbf{A})\mathbf{b}\|_{2}^{2}=\min_{\alpha_{0},\ldots,\alpha_{k-1}}\sum_{i=1}^{k+1}b_{i}^{2}\big{(}f(\lambda_{i})-p(\lambda_{i})\big{)}^{2}=\sum_{i=1}^{k+1}b_{i}^{2}\epsilon^{2}=\epsilon^{2}.$
Taking the square root of both sides gives the result.
A.2 Proof of Lemma 1.4
Proof A.3.
Let $\mathbf{V}$ be the matrix whose $i$th column is $\mathbf{v}_{i}$. Let $w_{i}=\mathbf{v}_{i}^{\mathsf{T}}\mathbf{b}$ and $\mathbf{w}=\mathbf{V}^{\mathsf{T}}\mathbf{b}$ be the vector whose $i$th entry is $w_{i}$. Then,
$\|f(\mathbf{A})\mathbf{b}-p(\mathbf{A})\mathbf{b}\|_{2}^{2}=\sum_{i=1}^{d}(f(\lambda_{i})-p(\lambda_{i}))^{2}w_{i}^{2}\geq\left(\min_{j}w_{j}^{2}\right)\sum_{i=1}^{d}(f(\lambda_{i})-p(\lambda_{i}))^{2}.$
Clearly $\sum_{i=1}^{d}(f(\lambda_{i})-p(\lambda_{i}))^{2}\geq\max_{x\in\Lambda}|f(x)-p(x)|^{2}$. Taking the square root of both sides and substituting into Definition 1.2 proves the result.
For completeness, we now show that if $\mathbf{b}\sim\mathcal{N}(\mathbf{0},\mathbf{I}_{d})$, then with high probability,
$\frac{\|\mathbf{b}\|_{2}}{\min_{j}|\mathbf{v}_{j}^{\mathsf{T}}\mathbf{b}|}=O(d^{3/2}).$
First, with high probability, $\|\mathbf{b}\|_{2}=\Theta(\sqrt{d})$; see, for example, [Ver18, 3.1.2]. Second, we show that with high probability $|\mathbf{v}_{j}^{\mathsf{T}}\mathbf{b}|=\Omega(1/d)$ for all $j$. Since $\mathbf{V}$ is orthonormal, $\mathbf{w}:=\mathbf{V}^{\mathsf{T}}\mathbf{b}\sim\mathcal{N}(\mathbf{0},\mathbf{I}_{d})$. By anti-concentration of the normal distribution, the probability that $|w_{j}|>\epsilon$ is at least $1-0.4\epsilon$ for any $\epsilon$. By a union bound, this holds simultaneously for all $j$ with probability at least $1-0.4\epsilon d$. Setting $\epsilon=\Theta(1/d)$ finishes the argument. Finally, by another union bound, our bounds on the numerator and on the denominator hold simultaneously with high probability.
A.3 Lanczos-FA on indefinite systems
The MINRES iterates are defined as
$\hat{\mathbf{y}}_{k}:=\operatornamewithlimits{argmin}_{\mathbf{y}\in\mathcal{K}_{k}(\mathbf{A},\mathbf{b})}\|\mathbf{b}-\mathbf{A}\mathbf{y}\|_{2}=\operatornamewithlimits{argmin}_{\mathbf{y}\in\mathcal{K}_{k}(\mathbf{A},\mathbf{b})}\|\mathbf{A}^{-1}\mathbf{b}-\mathbf{y}\|_{\mathbf{A}^{2}}.$
Define the residual vectors
$\mathbf{r}_{k}:=\mathbf{b}-\mathbf{A}\mathsf{lan}_{k}(1/x;\mathbf{A},\mathbf{b}),\qquad\mathbf{r}_{k}^{\mathrm{M}}:=\mathbf{b}-\mathbf{A}\hat{\mathbf{y}}_{k},$
and note that the MINRES residual norms are non-increasing due to the optimality of the MINRES iterates. In [CG96], it is shown that the CG residual norms are near the MINRES residual norms at iterations where MINRES makes good progress. More precisely, the algorithms are related by
$\|\mathbf{r}_{k}\|_{2}=\frac{\|\mathbf{r}_{k}^{\mathrm{M}}\|_{2}}{\sqrt{1-\left(\|\mathbf{r}_{k}^{\mathrm{M}}\|_{2}/\|\mathbf{r}_{k-1}^{\mathrm{M}}\|_{2}\right)^{2}}}.$
In [CGMM22, Appendix A], this result was used to show that Lanczos-FA has an effective geometric convergence rate. The following theorem, which we stated as Eq. 7, represents a strengthening of [CGMM22, Appendix A]. To the best of our knowledge, the following result is new.
Theorem A.4.
For every $k>0$, there exists $k^{*}\leq k$ such that
${\|\mathbf{b}-\mathbf{A}\mathsf{lan}_{k^{*}}(1/x;\mathbf{A},\mathbf{b})\|_{2}}\leq\left(\mathrm{e}\sqrt{k}+\frac{1}{\sqrt{k}}\right)\min_{\deg(p)<k}{\|\mathbf{b}-\mathbf{A}p(\mathbf{A})\mathbf{b}\|_{2}}.$
Switching from the 2-norm of the residual ($\mathbf{A}^{2}$-norm of the error) to the 2-norm of the error gives the bound stated in Eq. 7.
Proof A.5.
Fix $k>0$. If $\sqrt{\mathrm{e}}\|\mathbf{r}_{k}^{\mathrm{M}}\|_{2}>\|\mathbf{r}_{0}^{\mathrm{M}}\|_{2}=\|\mathbf{b}\|_{2}$, then the theorem holds with $k^{*}=0$ as $\sqrt{\mathrm{e}}\leq(\mathrm{e}\sqrt{k}+1/\sqrt{k})$. Assume $\sqrt{\mathrm{e}}\|\mathbf{r}_{k}^{\mathrm{M}}\|_{2}\leq\|\mathbf{r}_{0}^{\mathrm{M}}\|_{2}=\|\mathbf{b}\|_{2}$. Then there exists an integer $k^{\prime}\in[0,k)$ such that
$\|\mathbf{r}_{k^{\prime}}^{\mathrm{M}}\|_{2}\geq\sqrt{\mathrm{e}}\|\mathbf{r}_{k}^{\mathrm{M}}\|_{2}\geq\|\mathbf{r}_{k^{\prime}+1}^{\mathrm{M}}\|_{2}.$
This implies that
$\frac{1}{\sqrt{\mathrm{e}}}\geq\frac{\|\mathbf{r}_{k}^{\mathrm{M}}\|_{2}}{\|\mathbf{r}_{k^{\prime}}^{\mathrm{M}}\|_{2}}=\frac{\|\mathbf{r}_{k^{\prime}+1}^{\mathrm{M}}\|_{2}}{\|\mathbf{r}_{k^{\prime}}^{\mathrm{M}}\|_{2}}\cdot\frac{\|\mathbf{r}_{k^{\prime}+2}^{\mathrm{M}}\|_{2}}{\|\mathbf{r}_{k^{\prime}+1}^{\mathrm{M}}\|_{2}}\>\cdots\>\frac{\|\mathbf{r}_{k}^{\mathrm{M}}\|_{2}}{\|\mathbf{r}_{k-1}^{\mathrm{M}}\|_{2}}.$
Thus, there exists an integer $k^{*}\in(k^{\prime},k]$ such that
$\frac{\|\mathbf{r}_{k^{*}}^{\mathrm{M}}\|_{2}}{\|\mathbf{r}_{k^{*}-1}^{\mathrm{M}}\|_{2}}\leq\left(\frac{1}{\sqrt{\mathrm{e}}}\right)^{1/(k-k^{\prime})}\leq\left(\frac{1}{\sqrt{\mathrm{e}}}\right)^{1/k}.$
Since the MINRES residual norms are non-increasing, $\|\mathbf{r}_{k^{*}}^{\mathrm{M}}\|_{2}\leq\|\mathbf{r}_{k^{\prime}+1}^{\mathrm{M}}\|_{2}\leq\sqrt{\mathrm{e}}\|\mathbf{r}_{k}^{\mathrm{M}}\|_{2}$. Thus,
$\|\mathbf{r}_{k^{*}}\|_{2}=\frac{\|\mathbf{r}_{k^{*}}^{\mathrm{M}}\|_{2}}{\sqrt{1-\left(\|\mathbf{r}_{k^{*}}^{\mathrm{M}}\|_{2}/\|\mathbf{r}_{k^{*}-1}^{\mathrm{M}}\|_{2}\right)^{2}}}\leq\frac{\sqrt{\mathrm{e}}\|\mathbf{r}_{k}^{\mathrm{M}}\|_{2}}{\sqrt{1-(1/\sqrt{\mathrm{e}})^{2/k}}}.$
It can be verified that
$\mathrm{e}\sqrt{k}\leq\frac{\mathrm{e}}{\sqrt{1-(1/\mathrm{e})^{1/k}}}\leq\mathrm{e}\sqrt{k}+\frac{1}{\sqrt{k}}.$
Thus, we find that
$\|\mathbf{b}-\mathbf{A}\mathsf{lan}_{k^{*}}(1/x;\mathbf{A},\mathbf{b})\|_{2}=\|\mathbf{r}_{k^{*}}\|_{2}$ $\leq\left(\mathrm{e}\sqrt{k}+\frac{1}{\sqrt{k}}\right)\|\mathbf{r}_{k}^{\textup{M}}\|_{2}.$
The result follows from the optimality of the MINRES iterates.